What is the angular momentum of a 2.9-kg uniform c
Answer #1

here,

mass , m = 2.9 kg

radius , r = 29 cm = 0.29 m

angular speed , w = 1400 rpm

w = 146.5 rad/s

the moment of inertia , I = 0.5 * m * r^2

I = 0.5 * 2.9 * 0.29^2 = 0.12 kg.m^2

the angular momentum , L = I * w

L = 0.12 * 146.5 kg.m^2 /s

L = 17.86 kg.m^2/s

the angular momentum is 17.86 kg.m^2/s